OpenStudy (anonymous):

Determine if the sequence 1, 2, 3, 5, 8, 13... is arithmetic, geometric, or niether. A. Arithmetic B. Geometric C. Niether

OpenStudy (amistre64):

is 1/2 the same as 2/3 ?

OpenStudy (anonymous):

No

OpenStudy (amistre64):

then which one can we rule out?

OpenStudy (anonymous):

1/2

OpenStudy (amistre64):

1/2 isnt an option :/ 1/2 and 2/3 tell us something about a ratio between terms

OpenStudy (amistre64):

does: 1-2 = 2-3 = 3-5 ?

OpenStudy (anonymous):

Nope i get -1

OpenStudy (amistre64):

1, 2, 3, 5, 8, 13 IF: 1/2 =2/3= 3/5 = 5/8 = 8/13 ; then the series would have a common ratio IF: 1-2 = 2-3 = 3-5 = 5-8 = 8-13 ; then we would have a common difference

OpenStudy (amistre64):

a common ratio is geometric; a common difference is arithmetic ... otherwise its none of them

OpenStudy (anonymous):

so its neither